Roxie is picking out some movies to rent, and she is primarily interested in horror films and foreign films. She has narrowed down her selections to 20 horror films and 8 foreign films.

Step 1 of 2 :  

How many different combinations of 3 movies can she rent?
Answer #1

There are total 20 + 8 = 28 films,

3 films are selected from 28 films by 28C3 ways,

28C3 = (28! / [ (28-3)! * 3! ]

= 28 ! / [ (25 ! * 3! ]

= 3276
Answer #2

To calculate the number of different combinations of 3 movies Roxie can rent, we can use the concept of combinations.

In mathematics, the number of combinations of selecting r objects from a set of n objects, without regard to the order, is given by the formula:

C(n, r) = n! / (r! * (n - r)!)

Where: C(n, r) represents the number of combinations of selecting r objects from a set of n objects. n! (n factorial) is the product of all positive integers from 1 to n. r! (r factorial) is the product of all positive integers from 1 to r. (n - r)! is the product of all positive integers from 1 to (n - r).

In Roxie's case: n = Total number of movies available = 20 (horror films) + 8 (foreign films) = 28 r = Number of movies she wants to rent = 3

Let's calculate the number of combinations:

C(28, 3) = 28! / (3! * (28 - 3)!) C(28, 3) = (28 * 27 * 26) / (3 * 2 * 1) C(28, 3) = 3276

So, Roxie can rent 3 movies in 3,276 different combinations from the available selection of horror films and foreign films.

answered by: Hydra Master